# Keyframer: Empowering Animation Design using Large Language Models

### v1

@@ Abstract.
Large language models (LLMs) have the potential to impact a wide range of creative domains, but the application of LLMs to animation is underexplored and presents novel challenges such as how users might effectively describe motion in natural language. In this paper, we present Keyframer, a design tool for animating static images (SVGs) with natural language. Informed by interviews with professional animation designers and engineers, Keyframer supports exploration and refinement of animations through the combination of prompting and direct editing of generated output. The system also enables users to request design variants, supporting comparison and ideation. Through a user study with 13 participants, we contribute a characterization of user prompting strategies, including a taxonomy of semantic prompt types for describing motion and a ‘decomposed’ prompting style where users continually adapt their goals in response to generated output. We share how direct editing along with prompting enables iteration beyond one-shot prompting interfaces common in generative tools today. Through this work, we propose how LLMs might empower a range of audiences to engage with animation creation.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Human-centered computing Interactive systems and tools
@@ 1. Introduction
There is great potential for large languages models (LLMs) to influence and support creative work across the design process, from concept development to refinement to execution. Over the past year alone, researchers have developed LLM-powered design tools in a range of domains such as visual design (Chiou et al., 2023; Sanchez, 2023; Choi et al., 2023), creative writing (Gero et al., 2023; Suh et al., 2023), and 3D modeling (Gmeiner et al., 2023; Liu et al., 2023). The use of descriptive natural language prompts has the potential to reduce the technical expertise needed to create design artifacts and provide novices with educational opportunities to improve their design skills (Jonsson and Tholander, 2022). Yet, while prompting strategies for text-to-image generators have been well studied in recent work (Chiou et al., 2023; Sanchez, 2023; Liu and Chilton, 2022; Kulkarni et al., 2023), the transferability of these strategies to new domains is uncertain, and research is needed to understand domain-specific needs.
In this paper, we apply LLMs to the less explored domain of animation design. Using LLMs in this domain may be especially fruitful as creating animations requires many different types of technical skills, from applying motion design principles for compelling visual communication, to executing and implementing animations in code in production (such as for advertising, games, and user interfaces); as a result, animation work often involves varied stakeholders like motion designers, technical artists, and software engineers. While one-shot prompting interfaces are common in commercial text-to-image systems like Dall·E [^1] https://openai.com/dall-e-3 and Midjourney [^2] https://www.midjourney.com/, we argue that animations require a more complex set of user considerations, such as timing and coordination, that are difficult to fully specify in a single prompt—thus, alternative approaches that enable users to iteratively construct and refine generated designs may be needed especially for animations.
@@
We combined emerging design principles for language-based prompting of design artifacts with code-generation capabilities of LLMs to build a new AI-powered animation tool called Keyframer. With Keyframer, users can create animated illustrations from static 2D images via natural language prompting. Using GPT-4  [^3] https://openai.com/research/gpt-4, Keyframer generates CSS animation code to animate an input Scalable Vector Graphic (SVG). To support user refinement of generated designs, Keyframer has multiple editor types for users to directly edit generated animations. Additionally, users can iterate on their designs through sequential prompting and request variants from the LLM to ideate on new design directions. Through these features, Keyframer accommodates users exploring and adapting their design goals as they iteratively construct an animation through combined prompting and editing actions.
@@
To inform the development of Keyframer, we first conducted formative interviews with 9 professional designers and engineers who work on animations, such as character design and UI design of mobile applications. Through these interviews, we aimed to answer the research question,
@@
Through our analysis of user prompting strategies, we contribute a taxonomy of semantic prompting styles users employed to describe motion with natural language, showing how Keyframer empowered both novices and experts to create animations while focusing on high level design goals. We describe how users predominantly built up their designs through a sequential prompting style we call ‘decomposed’ prompting. With decomposed prompting, users iteratively develop their ideas through incremental prompts that animate individual elements within a scene. This workflow can allow users to adapt their goals continually in response to generated output. By providing multiple pathways for iterating on designs (through a prompting interface and editors for modify generated animation properties), Keyframer allows users to maintain creative control throughout their design process. Further, we share how unexpected LLM output can be beneficial both in exploration and refinement stages of the design process. Our work makes the following contributions:
@@
    Keyframer, an LLM-powered application for generating animations from static images.
@@
    A taxonomy of prompting strategies users employed to describe animations with natural language.
  3. (3)
    Insights into supporting design iteration through combined prompting and editing interfaces.
@@ 2.1. Generative AI in Design
The introduction of LLMs like ChatGPT [^4] https://openai.com/chatgpt have facilitated an unprecedented rise in commercial and research efforts to explore their application to design fields. Muller et al. argue that generative AI introduces new challenges to the field of HCI “due to the serendipitous and uncertain nature of the design space,” with numerous open questions about how to design user experiences that can effectively facilitate creative work (Muller et al., 2022). The use of natural language input presents opportunities to lower the barrier to entry for engaging in creative practice, while the interface paradigms of co-creating with AI are being proposed in domains such as graphic design (Chiou et al., 2023; Sanchez, 2023; Feng et al., 2023; Liu and Chilton, 2022), software development (Barke et al., 2023; Vaithilingam et al., 2022), creative computing (Angert et al., 2023; Jonsson and Tholander, 2022), UI design (Petridis et al., 2023), writing (Gero et al., 2023; Suh et al., 2023), 3D CAD (Liu et al., 2023; Gmeiner et al., 2023), and music (Louie et al., 2020).
A common challenge in natural-language-based generative AI tools is developing effective prompting strategies for steering generated output. Today, prompting tends to require much trial and error, with a lack of meaningful controls for end users (Liu and Chilton, 2022). Current efforts to define prompting strategies largely center on text-to-image generators that tend to employ one-shot approaches, where the user’s only lever to edit the image is via text prompting (though researchers are also exploring multi-modal input with images and painting interfaces for editing selected regions (Dang et al., 2023; Liu and Chilton, 2022; Singh et al., 2022; Artbreeder, 2022; Elharrouss et al., 2020)). Several prompt taxonomies have been proposed (Oppenlaender, 2023; Sanchez, 2023), with generative art communities using modifiers specifying artistic style (e.g., ‘Cubism’) and quality (e.g., ‘award winning’), along with keywords to spur surprising output (what Oppenlaender refers to as ‘magic terms’ (Oppenlaender, 2023)). Similarly, Chiou et al distinguish between ‘operational’ keywords that specify concrete reference terms and ‘conceptual’ keywords using abstract modifiers that are more likely to lead to unexpected results (Chiou et al., 2023). Emerging research on LLM-generated scalable vector graphics (SVGs) suggests a promising path towards recombining high-quality, designer-created assets for consistent visual style (Miriam Jacob et al., 2024).
Early work on applying LLMs to create animated content has appeared in the form of creative coding support for p5js [^5] https://p5js.org/(Angert et al., 2023) and video generation with tools like RunwayML [^6] https://runwayml.com/. However, research exploring generative AI for creating animations from existing image assets has been underexplored; this process is most closely aligned with current professional practices, where animations are created from assets created by visual designers. Research is needed to determine whether animations may require alternative prompting strategies compared to text-to-image generation, or even to determine how effective LLMs may be at generating animations altogether.
To mitigate the trial-and-error workflow common to prompting generative AI tools, several prototype tools have sought to visualize related keywords to help users better explore their options (Feng et al., 2023; Suh et al., 2023; Liu et al., 2023; Brade et al., 2023). Commercial image generators like Dall-E  [^7] https://openai.com/dall-e-3 and Midjourney [^8] https://www.midjourney.com/ often present users with polished visual output with only a handful of keywords; yet, this may lead users to narrow in on one design direction before fully exploring alternatives (Suh et al., 2023). Presenting users with multiple options has been found to support users overcoming creative blocks in their process (Choi et al., 2023; Angert et al., 2023), spur new ideas (Tholander and Jonsson, 2023; Suh et al., 2023), and verify the quality of generated output (Barke et al., 2023). In the context of code generation, other related work has similarly distinguished between two use cases for LLMs: supporting ‘exploration,’ when a user is not sure of what they want to create yet, and ‘acceleration,’ where a task is well defined and the user wants the LLM to help them get to a solution faster (Barke et al., 2023; Angert et al., 2023; Jonsson and Tholander, 2022). In this work, we are interested in how prompting for animations might compare to the existing literature on prompting in areas like image generation and prose. Further, we consider how the integration of design variants might help support ideation processes and ultimately promote design iteration.
@@ 2.2. Iteration in Design Prototyping
Iteration is a fundamental and crucial aspect of the design process. Typically consisting of repeatedly exploring new ideas and refining designs, iteration can foster alternative perspectives (Dow et al., 2011), reflection (Hartmann et al., 2006; Adams et al., 2003), and stakeholder feedback (Dow et al., 2011; Yen et al., 2017), aiding designers in identifying challenges and uncovering new directions (Buxton and Sniderman, 1980; Kelly, 2002; Adams, 2002; Camburn et al., 2017; Dunlap et al., 2014). Iteration is essential to complex designs where exploring different alternatives can reveal potential issues and opportunities. Animation design exemplifies a domain that greatly benefits from iteration due to its multiple stages and stakeholders. For example, film animations involves the integration of story, script, storyboarding, media design, and visual narratives (Tsai-Yun et al., 2013), while animations for games can involve character rigging, lighting, and special effects (Heller and Womack, 2011). Creating animations often requires collaboration among diverse stakeholders and iteration to refine ideas and ensure alignment across all elements of the project. Our work with Keyframer focuses on computer-based animation, specifically the creation of web-based animations, which often involves graphic designers and front-end engineers working together to create production-ready animations such as animated transitions, loading animations, and data visualizations (Drasner, 2017; Head, 2016).
@@
Iteration involves both exploration and refinement, as illustrated by the classic “Double Diamond” model in design (Council, 2004). Exploration typically entails generating ideas and multiple variants of a design solution, where designers experiment with different styles and design languages. Instead of committing to a single design concept at the early stage of the design process, creating design variants can help designers avoid the pitfall of fixation (Jansson and Smith, 1991), discover potentially valuable directions (Cross, 2004, 2006), and generate a more diverse array of ideas (Buxton, 2010). An effective practice of creating design variants is parallel prototyping, which involves creating multiple design variants in parallel before making further improvements on any of the variants (Camburn et al., 2017). By generating and comparing varied prototypes in parallel, designers were able to create higher quality and more diverse design solutions and feel more confident in the design process than working on single designs (Dow et al., 2010, 2009a; Neeley Jr et al., 2013). Refinement is where designers go deep on improving a single design to match their design goals (Ball et al., 1997). It has been proven in multiple studies that rapid refinement on existing designs can enhance design outcomes (Dow et al., 2009b). In this paper, we explore how the emerging technology of large language models can assist exploration and refinement in animation design.
@@ 3. Formative Study
We recruited animation designers, developers, and prototypers via Slack channels dedicated to animation and front-end engineering at a large technology company. We invited participants who have ‘experience designing and/or developing animations as part of your work’ to fill out a short survey with background questions such as their job title and what tools they use to create their animations. Selected participants received a \$12 meal voucher for their time.
During our 45-minute, semi-structured, video conference-based interviews, we asked participants about the types of animations they have worked on in the past and then asked each participant to walk us through a specific example where they created an animation using screen sharing. We then asked about their perspective of the role of generative AI in design tools. Finally, we probed the participants with a brief demo of an early prototype of Keyframer that takes in natural language prompts to generate animation for a simple SVG. We used this demo as a probe to ask for feedback about potential benefits or limitations of LLM-based animation tools.
To analyze our interviews, two researchers edited and reviewed the transcriptions automatically generated by the video conferencing tool and applied thematic analysis (Braun and Clarke, 2006) to inductively identify a set of painpoints and opportunities across different stages of the animation creation process. We also categorized different opportunities for generative AI support shared by participants when they were describing their experience with existing generative AI design tools and reacting to the concept of LLM-based animation tools that we showed in the demo. We conducted the interviews in sets of three, analyzing and discussing emergent themes in weekly meetings. We continued the interviews until we reached saturation, with no new emergent themes, which happened after 9 total interviews.
@@ 3.1. Formative Interview Results
From a total of 27 survey respondents, we selected and interviewed 9 participants (2 female, 7 male) representing a range of job titles such as designer, technical R&D artist, and front-end developer. Their animation work includes user interface animations, advertising, instructional documentation, character animation, and interactive data visualization with between 3–22 years of professional experience (details about our participants can be found in Table A.1).  [^9] We use FP to refer to formative interview participants and EP to refer to evaluative study participants in Section 6 While a full discussion of challenges identified throughout the entire animation lifecycle is out of scope for this paper, we highlight a few opportunities especially pertinent to the role generative AI might play in animation design tooling.
@@
Challenge: Translating design to engineering implementation. Transitioning from design to engineering implementation (commonly referred to as “handoff”) is time-consuming and can ultimately take away from other stages of the animation design process. Translation happens both because design is a collaborative process between designers and engineers, and because of differences in performance of animations across tools (e.g., production designs may appear differently than design mockups due to latency or rendering discrepancies): “In animation software, I can do anything, but real product might have limitations” (FP3). Handoff often manifests in the creation of design specs or the development of custom tooling for translating animation properties in software like After Effects into production code (Swift, JavaScript, etc). Translation work, especially when it can involve several rounds of feedback and adjustment between prototyping and production, takes time away from design iteration: “If I do one animation and try it and then it takes me 3 hours till I get to see it on the product, I don’t get many cycles [of iteration] and then the animation quality drops significantly” (FP3). For these reasons, there was a general preference to get as close as possible to the final format the animation will ultimately appear in, which for the creators we interviewed, was either implemented in code on a webpage or in a mobile application. Participants like FP1 imagined how generative tools that generate both animation and the underlining code could mitigate this challenge: “It would help because the code is there, so it will also help an engineer understand [the design]… it’s a communication tool for both [engineering and design].”
Opportunity: Using AI to help generate an initial starting point. Animation creators saw potential in using AI to generate an initial starting point that they could then refine. Prototype engineer FP6 described his enthusiasm for tools like Midjourney: “It’s a great starting point for a lot of things, and it helps unblock that creative process a lot of the times of, like, hey, where do I start?” Front-end developers anticipated the usefulness of AI for writing code: “Eventually, it [AI design tools] is going to be very helpful — not to write all the code, but to help in certain parts of the code” (FP2).
Even with potential benefits of using AI, designers emphasized the importance of ultimately customizing the AI-generated output, rather than relying on the AI to produce final designs, as designer FP6 described: “If an [AI] could help alleviate and create some more complex animations, or at least start it for me so I don’t have to write every single line of code, that would save me so much time… I’m not writing the actual base code, I’m just truly putting that human touch to it.”
@@
Opportunity: Supporting design iteration. Another area animation creators saw potential for generative AI was in iterating on animations, particularly by helping them explore multiple options. Some participants like FP7 imagined LLMs serving more to refine designs rather than proposing completely new designs, suggesting using AI to apply design standards like accessible complementary colors or a dark mode from a light mode design. FP5 shared how it was rare for animators to create a completely new timing curve, so applying animation properties from references examples would be valuable. Designer FP1 shared how AI can be especially helpful in generating design variants quickly from an initial design: “I feel like this can be super helpful, especially for those things that need to be tweaked.”
@@
Limitations: Lack of Creative Control. A general concern about using generative AI was the risk of losing creative control, both in terms of reduced editing functionality along with a loss in creative satisfaction. For example, FP4 shared, “[When] AI is being used to create the entire work, it doesn’t scratch the itch that I think creative people have when they have that desire to create something. [With Midjourney] I feel like I’m pushing a button and something pops out.” Participant FP9 also shared that, “It’s the granularity of customization that I really struggle with. When you’re drawing, you control every stroke, and the nuance of it matters, and you control so little in AI by definition.” Designers also feared having decreased understanding of how to edit generated designs. Developer FP8 described this gap by stating that, “If you involve yourself deeply in all the details of the work, you have more freedom because you know exactly where things could break… [AI could] obfuscate some of these layers. And that could make it harder, if something goes wrong, to know where to look because that code was essentially not written by you.”

### v2

@@ Abstract.
Creating 2D animations is a complex, iterative process requiring continuous adjustments to movement, timing, and coordination of multiple elements within a scene. To support designers of varying levels of experience with animation design and implementation, we developed Keyframer, a design tool that generates animation code in response to natural language prompts, enabling users to preview rendered animations inline and edit them directly through provided editors. Through a user study with 13 novices and experts in animation design and programming, we contribute 1) a categorization of semantic prompt types for describing motion and identification of a ‘decomposed’ prompting style where users continually adapt their goals in response to generated output; and 2) design insights on supporting iterative refinement of animations through the combination of direct editing and natural language interfaces.
@@
$c$s: Human-centered computing Interactive systems and tools
@@ 1. Introduction
Animation design lives at the intersection of visual design and code (Reynolds, 1982). In application domains such as advertising, games, and user interfaces, creating animations requires a range of expertise in areas such as motion design, technical art, and front-end engineering. Like other creative work, animation design is inherently iterative, involving the concurrent specification and refinement of properties like movement and timing for multiple elements in a scene, which necessitates continuous adjustment and translation between artistic concepts and technical implementation. As a result, animation design, and the engineering needed to implement animations into applications, remains a challenging task for creators of all skill levels.
Given promising early results applying large language models to support designers in a range of artistic domains (such as visual design (Chiou et al., 2023; Sanchez, 2023; Choi et al., 2023), creative writing (Gero et al., 2023; Suh et al., 2024), and 3D modeling (Gmeiner et al., 2023; Liu et al., 2023)), along with the ability of LLMs to quickly generate code to support software development [^1] https://github.com/features/copilot, we considered the potential of LLMs to support the animation design and engineering process. The ability of LLMs to interpret and respond to natural language requests may reduce the technical barriers needed to create animations, while also offering opportunities to support novices with both design and programming education (Jonsson and Tholander, 2022). Yet, while use of natural language-based systems have proliferated in recent years (such as for text-to-image generation with tools like Dall·E [^2] https://openai.com/dall-e-3 and Midjourney [^3] https://www.midjourney.com/ or code generation with GitHub Copilot [^4] https://github.com/features/copilot), these existing commercial tools are not designed for the specific needs of animation creation because they lack support for the highly-iterative process of continuous refinement and visualization of animation designs. During this process, small changes to animation properties require a tight feedback loop of design, preview, and editing which is not currently supported with AI tools that require time to regenerate for every iteration and new prompt.
Further, while prompting strategies for text-to-image generators have been well studied in recent work (Chiou et al., 2023; Sanchez, 2023; Liu and Chilton, 2022; Kulkarni et al., 2023), the transferability of these strategies to animation design is uncertain, as specifying how multiple items should animate together can be especially difficult to describe and plan for upfront. Therefore, to support animation design, we need alternative approaches to existing LLM-based design and programming tools, as well as deeper research insights into how animation designers might communicate their ideas and iteratively develop their work with the help of LLMs.
@@
To explore design opportunities for LLM-supported animation design, we created an AI-powered animation tool called Keyframer. With Keyframer, users can create animated illustrations from static 2D images using a combination of prompting and direct editing of animation properties. Once a user provides an input scalable vector graphic (SVG) along with a natural language prompt, Keyframer uses GPT-4555https://openai.com/research/gpt-4 to generate CSS animation code to animate the image. Because the generated animation is created using code, users can then directly edit animation properties with provided editors available in Keyframer. They can continually iterate on their designs through additional prompts or direct edits, and they can also request variants from the LLM to ideate on new design directions. Through these features, Keyframer accommodates users exploring and adapting their design goals as they iteratively construct an animation through combined prompting and editing actions, supporting the tight feedback loop needed for animation.
@@
To inform the development of Keyframer, we first conducted formative interviews with 9 professional designers and engineers who work on animations. Through these interviews, we aimed to answer the research question,
@@
Through our analysis of user prompting strategies, we contribute a categorization of semantic prompting styles users employed to describe motion with natural language, showing how Keyframer empowered both novices and experts to create animations while focusing on high level design goals. We also describe how Keyframer enables users to predominantly build up their designs through sequential ‘decomposed’ prompting, where users iteratively develop their ideas through incremental prompts that animate individual elements within a scene, as guided by generated output.
In addition, through providing an interface that enables users to request animation designs through natural language prompting, provides direct previews of animations alongside generated code, and enables direct edits to generated output, we show how Keyframer allows for design exploration and iteration throughout the 2D animation design process. Our work makes the following contributions:
@@
    Keyframer, an open-sourced [^6] https://github.com/apple/ml-keyframer LLM-powered application for designing animations from static images.
@@
    Insights into how iterative design practices with 2D animation were supported by Keyframer, drawn from an empirical exploratory study with 13 users
@@ 2.1. Generative AI in Design
The introduction of LLMs has facilitated an unprecedented rise in commercial and research efforts to explore their application to design fields. Muller et al. argue that generative AI introduces new challenges to the field of HCI “due to the serendipitous and uncertain nature of the design space,” with numerous open questions about how to design user experiences that can effectively facilitate creative work (Muller et al., 2022). Interface paradigms of co-creating with AI are being proposed in domains such as graphic design (Chiou et al., 2023; Sanchez, 2023; Feng et al., 2023; Liu and Chilton, 2022), software development (Barke et al., 2023; Vaithilingam et al., 2022), creative computing (Angert et al., 2023; Jonsson and Tholander, 2022), UI design (Petridis et al., 2023), writing (Gero et al., 2023; Suh et al., 2024), 3D CAD (Liu et al., 2023; Gmeiner et al., 2023), and music (Louie et al., 2020).
One common challenge with using natural-language-based AI design tools is learning to develop effective prompting strategies to steer generated output. The typical user journey for using generative tools today begins with crafting the overall structure of a prompt, followed by user assessment of generated designs and continuous refinement of prompts for better results (Mahdavi Goloujeh et al., 2024; Xie et al., 2023). Several user prompt taxonomies have been proposed in the domain of text-to-image generators (Oppenlaender, 2023; Sanchez, 2023), with artists frequently using modifiers specifying artistic style (e.g., ‘Cubism’) and quality (e.g., ‘award winning’), along with keywords to spur surprising output (what Oppenlaender refers to as ‘magic terms’ (Oppenlaender, 2023)). Similarly, Chiou et al distinguish between ‘operational’ keywords that specify concrete reference terms and ‘conceptual’ keywords using abstract modifiers that are more likely to lead to unexpected results (Chiou et al., 2023). We posit that animations involve complex and unique user considerations compared to images, including timing and coordination, that needs to be studied to best understand effective prompting guidelines for creating animated content.
Early work exploring the creation of animated content from natural language prompts include Spellburst, which provides creative coding support for p5js [^7] https://p5js.org/(Angert et al., 2023) and video generation with tools like RunwayML [^8] https://runwayml.com/ and OpenAI’s Sora [^9] https://openai.com/index/sora/. Yet, research exploring generative AI for creating animations from existing image assets has been underexplored; this process is most closely aligned with current professional practices, where animations are built from assets created by visual designers to ensure high quality and consistency. Emerging research on LLM-generated scalable vector graphics (SVGs) suggests a promising path (Miriam Jacob et al., 2024) but has been only applied to static images. Our system was designed and studies were conducted in late 2023, when, to the best of our knowledge, it was the first tool to apply LLMs to SVG-based animations. Since then, projects such as LogoMotion (Liu et al., 2024) and MoVer (Ma and Agrawala, 2025) have considered predominantly automated approaches for creating animated logos and text. However, recent work has demonstrated how the use of LLMs more broadly might contribute to homogenization of creative output or a reduction in critical thinking (Lee et al., 2025), especially when LLMs present ‘complete-seeming’ ideas (Anderson et al., 2024; Bommasani et al., 2022), which we argue automated approaches to animation lean towards. Thus, by focusing instead on how creators can specify animation goals from scratch (rather than in response to automated suggestions) our work points a way towards offering users more flexibility in crafting their creative vision.
@@ 2.2. Iteration in Design Prototyping
Iteration is a fundamental and crucial aspect of design process. Typically consisting of repeatedly exploring new ideas and refining designs, iteration can foster alternative perspectives (Dow et al., 2011), reflection (Hartmann et al., 2006; Adams et al., 2003), and stakeholder feedback (Dow et al., 2011; Yen et al., 2017), aiding designers in identifying challenges and uncovering new directions (Buxton and Sniderman, 1980; Kelly, 2002; Adams, 2002; Camburn et al., 2017; Dunlap et al., 2014). Iteration is essential to complex designs where exploring different alternatives can reveal potential issues and opportunities. Animation design exemplifies a domain that greatly benefits from iteration due to its multiple stages and stakeholders. For example, film animations involves the integration of story, script, storyboarding, media design, and visual narratives (Tsai-Yun et al., 2013), while animations for games can involve character rigging, lighting, and special effects (Heller and Womack, 2011). Creating animations often requires collaboration among diverse stakeholders and iteration to refine ideas and ensure alignment across all elements of the project. Our work with Keyframer focuses on computer-based animation, specifically the creation of web-based animations, which often involves graphic designers and front-end engineers working together to create production-ready animations such as animated transitions, loading animations, and data visualizations (Drasner, 2017; Head, 2016).
@@
Iteration involves both exploration and refinement, as illustrated by the classic “Double Diamond” model in design (Council, 2004). Exploration typically entails generating ideas and multiple variants of a design solution, where designers experiment with different styles and design languages. Instead of committing to a single design concept at the early stage of the design process, creating design variants can help designers avoid the pitfall of fixation (Jansson and Smith, 1991), discover potentially valuable directions (Cross, 2004, 2006), and generate a more diverse array of ideas (Buxton, 2010). An effective practice of creating design variants is parallel prototyping, which involves creating multiple design variants in parallel before making further improvements on any of the variants (Camburn et al., 2017). By generating and comparing varied prototypes in parallel, designers were able to create higher quality and more diverse design solutions and feel more confident in the design process than working on single designs (Dow et al., 2010, 2009a; Neeley Jr et al., 2013). Refinement is where designers go deep on improving a single design to match their design goals (Ball et al., 1997). It has been demonstrated in multiple studies that rapid refinement on existing designs can enhance design outcomes (Dow et al., 2009b).
Despite the importance of iteration in design, commercial image generators like Dall-E  [^10] https://openai.com/dall-e-3 and Midjourney [^11] https://www.midjourney.com/ are limited in supporting users with meaningfully iterating on their ideas. These tools often present users with polished visual output, which may lead users to narrow in on one design direction before fully exploring alternatives (Suh et al., 2024). Emerging research has started to explore ways to support iteration in design using generative AIs. For example, (Zhou et al., 2024) presents a rule-based AI chat system for user interface design that facilitates the non-linear design process and generates alternatives for users to remix and refine. Presenting users with multiple options has been found to support users overcoming creative blocks in their process (Choi et al., 2023; Angert et al., 2023), spur new ideas (Tholander and Jonsson, 2023; Suh et al., 2024), and verify the quality of generated output (Barke et al., 2023). In the context of code generation, other related work has similarly distinguished between two use cases for LLMs: supporting ‘exploration,’ when a user is not sure of what they want to create yet, and ‘acceleration,’ where a task is well defined and the user wants the LLM to help them get to a solution faster (Barke et al., 2023; Angert et al., 2023; Jonsson and Tholander, 2022). While iteration and exploration are also crucial in animation designs, less is known about how generative AI can assist exploration and refinement in this domain.
In summary, research is needed to determine whether animations may require alternative prompting approaches compared to text-to-image generation strategies, or even to determine how effective LLMs may be at generating animations altogether. Our work examines how natural language prompting and code generation-based approaches could support animation design. Further, we consider how the integration of previews of LLM-generated designs, design variants, and code editing mechanisms might support iteration for animation design.
@@ 3. Formative Study
We recruited professional animation designers, developers, and prototypers via Slack channels dedicated to animation and front-end engineering at a large technology company [^12] Both of our formative study (Section 3) and user study (Section 5) were reviewed by the human subject research review board internal to our company. Selected participants received a \$12 meal voucher for per 45-minute they spent in our studies..
During our 45-minute semi-structured interviews over video conference, we asked participants to share about their current animation design process and perspective of the role of generative AI for design. Finally, we probed the participants with a brief demo of an early prototype of Keyframer to probe for feedback about LLM-based animation tools, with the facilitator demonstrating the system’s ability to generate animations for a simple SVG using natural language (participants did not directly interact with Keyframer). Participants received a \$12 meal voucher for their time.
Two researchers reviewed the session transcripts and applied thematic analysis (Braun and Clarke, 2006) to inductively identify a set of painpoints and opportunities across different stages of the animation creation process, as well as opportunities for generative AI support. Our protocol was reviewed and approved by the research ethics committee at the large technology company.
@@ 3.1. Formative Interview Results
We interviewed 9 participants (2 female, 7 male) representing a range of job titles such as designer, technical R&D artist, and front-end developer. Their animation work includes user interface animations, advertising, instructional documentation, character animation, and interactive data visualization with between 3–22 years of professional experience (details about our participants can be found in Appendix A.1). [^13] We use FP to refer to formative interview participants and EP to refer to evaluative study participants in Section 6 While a full discussion of challenges identified throughout the entire animation lifecycle is out of scope for this paper, we highlight a few opportunities especially pertinent to the role generative AI might play in animation design tooling.
@@
Challenge: Translating design to engineering implementation. Transitioning from design to engineering implementation (commonly referred to as “handoff”) is time-consuming and can ultimately take away from other stages of the animation design process. Translation happens both because design is a collaborative process between designers and engineers, and because of differences in performance of animations across tools (e.g., production designs may appear differently than design mockups due to latency or rendering discrepancies). Translation work, especially when it can involve several rounds of feedback and adjustment between prototyping and production, takes time away from design iteration: “If I do one animation and try it and then it takes me 3 hours till I get to see it on the product, I don’t get many cycles [of iteration] and then the animation quality drops significantly” (FP3). Participants like FP1 imagined how generative tools that create both animation and the underlining code could mitigate this challenge: “It would help because the code is there, so it will also help an engineer understand [the design]… it’s a communication tool for both [engineering and design].”
@@
Opportunity: Supporting design iteration. Animation creators saw potential in using AI to generate a starting point they could then refine: “If an [AI] could help create some more complex animations, or at least start it for me so I don’t have to write every single line of code, that would save me so much time… I’m not writing the base code, I’m just truly putting that human touch to it” (FP6) Animation creators also saw potential for generative AI to support iterating on animations, particularly by helping them explore multiple options. Participant FP7 imagined LLMs being most useful for refining designs rather than proposing completely new designs, suggesting using AI to apply design standards like accessible complementary colors or a dark mode from a light mode design. FP5 described how it was rare for animators to create a completely new timing curve, so applying animation properties from references examples would be valuable. Designer FP1 shared how AI can be especially helpful in generating design variants quickly from an initial design, tweaking designs so multiple could be compared at once.
@@
Limitations: Lack of Creative Control. A general concern about using generative AI was the risk of losing creative control, both in terms of reduced editing functionality along with a loss in creative satisfaction. For example, FP4 shared, “[When] AI is being used to create the entire work, it doesn’t scratch the itch that I think creative people have when they have that desire to create something. [With Midjourney] I feel like I’m pushing a button and something pops out.” Participant FP9 also shared that, “It’s the granularity of customization that I really struggle with. When you’re drawing, you control every stroke, and the nuance of it matters, and you control so little in AI by definition.” Designers also feared having decreased understanding of how to edit generated designs: “If you involve yourself deeply in all the details of the work, you have more freedom because you know exactly where things could break… [AI could] obfuscate some of these layers. That could make it harder, if something goes wrong, to know where to look because that code was essentially not written by you” (FP8).